# Search for Quadruplet Scalars using Boosted Decision Trees at the LHC

### v1

@@ Abstract
Beyond the Standard Model scenarios introduce additional scalar and fermion multiplets, which influence neutrino mass generation mechanisms and yield distinctive collider signatures. This work focuses on a particular scenario involving a neutral fermion quintuplet and a scalar quadruplet. The study examines the production and decay of the scalar quadruplet components at the LHC, emphasizing how their decay patterns, fermiophobic versus fermiophilic, depend on mass differences and Yukawa couplings with the fermion multiplets. A detailed collider analysis targeting final states with at least four leptons and two jets is conducted. The study also incorporates Standard Model backgrounds, leveraging multivariate techniques via Boosted Decision Trees. Results indicate discovery potential for scalar masses around 600–700 GeV and exclusion sensitivity extending beyond 1 TeV, highlighting the model’s promising experimental signatures and its role in probing new physics at colliders.
@@ 1 Introduction
The Standard Model (SM), while spectacularly successful, is fundamentally incomplete. Its inability to explain phenomena such as neutrino masses, the nature of dark matter, and the hierarchy problem, among other issues, strongly motivates physics beyond the SM (BSM). Many sophisticated models are proposed to overcome the shortcoming of SM, for example, supersymmetry [Haber:1984rc], 5D holographic theories [Maldacena:1997re, Gubser:1998bc, Witten:1998qj, Randall:1999ee], GUT [Georgi:1974sy], Pati-Salam Model [Pati:1974yy], extra dimensions [Arkani-Hamed:1998jmv, Randall:1999ee], Composite Higgs Models [Kaplan:1983fs, Agashe:2004rs], Left-Right symmetric models [Mohapatra:1974gc, Senjanovic:1975rk] and twin Higgs models [Chacko:2005pe]. In many of these models, extra scalar degrees of freedom are either necessary or appear as a consequence of the fundamental theory.
These scalar extensions are often studied within models that are well-motivated phenomenologically such as two Higgs doublet Models (2HDM) [Branco:2011iw], Type II seesaw with scalar Triplets [Schechter:1980gr], Georgi-Machacek Model [Georgi:1985nv]. It is possible to achieve tiny neutrino masses by the seesaw mechanism. However, attempts to reconcile small neutrino masses with experimental accessibility results in a tension between the seesaw scale and the size of the couplings (or other related parameters) in these tree-level seesaw models [Boucenna:2014zba, Ding:2014nga]. A resolution to this is to suppress the tree-level contribution and let the loop-diagram be the leading contribution to the dimension-5 operator [Bonnet:2012kz, Ma:2006km]. An alternative approach is to raise the dimension of the effective operator, as studied in Ref. [Babu:2009aq, Kumericki:2012bh, Nomura:2017abu, Ghosh:2018drw]. In these type of models, large scalar multiplets appear naturally. Another method to lower the seesaw scale is by the method of “cascade seesaw mechanism” [Ding:2014nga] wherein the neutrino masses are generated at tree level via a combination of Type-I and Type-III seesaw mechanisms, including both scalar and fermion multiplets as the exotic particle content.
@@
We study the production and decays of the quadruplet scalar (comprising of neutral, singly charged, and doubly charged components) at the LHC. The decay modes of these quadruplet scalars depend critically on the mass difference between the scalar and fermion multiplets. The Yukawa coupling between the scalar and fermion multiplets also plays a major role in their decays. There are two different decay patterns wherein the scalar can either decay to gauge bosons directly or they first decay to quintuplet fermions which subsequently decay to SM particles. These two scenarios are named as fermiophobic and fermiophilic, respectively. We discuss the possible collider signatures in these two scenarios along with the existing limits from the collider studies.
The introduction of two distinct $SU(2)$ multiplets as the BSM particle content poses significant challenges for high-energy collider data. Focusing on the pair and associated production of charged scalars, our detailed analysis of signal processes establishes that the final state characterized by $\geq 4$ leptons and $2$ jets (fermiophilic scenario) provides sufficient sensitivity to probe both, the quartet scalar and the quintuplet fermion, at the LHC. We study the signals for scalars in the mass range 3001000300-1000 GeV. We analyze the most relevant SM background exhibiting overlapping kinematics, with focus on the key reconstructed observables that offer some discrimination power. The sensitivity in these channels is estimated by performing a multivariate analysis (MVA) using the Boosted Decision Tree (BDT) framework.
@@ 2 Model
This work is motivated by the models studied previously in [Kumericki:2012bh, Nomura:2017abu, Kumar:2019tat]. To keep the analysis minimal and transparent, we consider the simplest scenario where the BSM particle content comprises a quadruplet scalar ($\Phi_{4}=\left(\phi^{++},\phi^{+}_{2},\phi^{0},\phi^{-}_{1}\right)^{T}$) and a quintuplet fermion ($\Sigma_{R}=\left[\Sigma_{1}^{++},\Sigma_{1}^{+},{\Sigma^{0}},\Sigma_{2}^{-},\Sigma_{2}^{--}\right]_{R}^{T}$) only. We denote the scalar and the fermion masses by $M_{\Phi}$ and $M_{\Sigma}$, respectively  [^5] The mass of the scalar comes from the scalar potential $V=\mu^{2}H^{\dagger}H+M_{\Phi}^{2}\Phi_{4}^{\dagger}\Phi_{4}$..
The kinetic term of the quadruplet scalar, obtained by expanding the tensors, is,
@@
where $m$ is the weak isospin and $s_{W}~(c_{W})=\sin\theta_{W}~(\cos\theta_{W})$, $\theta_{W}$ being the Weinberg angle. From here, we obtain the interactions responsible for production of the scalars viz.,
@@
Here, the neutral component of ${\Phi}_{4}$ gets a non-zero VEV, $\frac{v_{4}}{\sqrt{2}}$, induced by virtue of its quartic coupling with the SM doublet Higgs. The VEV is constrained from the measurement of the $\rho$ parameter [Kumericki:2012bh, Yu:2015pwa, Nomura:2017abu].  [^6] These models are further constrained from muon (g2)(g-2) and lepton flavor violation experiments which is discussed in [Nomura:2017abu, Kumar:2019tat].. The decay of the scalars to the gauge bosons is governed by the following terms:
@@
It is important to note that decay of the $\Phi_{4}$ to gauge bosons depends upon its VEV, $v_{4}$.
The Yukawa interaction between the quadruplet scalar and the quintuplet fermion can be expressed as,
@@
This Yukawa interaction induces diagrams that lead to tiny neutrino masses [Nomura:2017abu, Kumar:2019tat]. The induced VEV allows sizeable Yukawa coupling for quintuplet masses $\leq 1$ TeV as neutrino mass, $m_{\nu}\propto v_{4}^{2}~y_{\nu}M_{\Sigma}^{-1}y_{\nu}^{T}$ [Nomura:2017abu]. For example, we find that for small $y_{\nu}$ ($\sim 0.0007$), $v_{4}\sim$ 1 GeV and for large $y_{\nu}$ ($\sim 1$), $v_{4}\sim 0.0002$ GeV. In order to generate non-zero masses for all three SM neutrinos, we need three copies of the quintuplets [Casas:2001sr, Sahdev:2024wie]. For the purpose of this study, we assume that $M(\Sigma_{2},~\Sigma_{3})>>M(\Sigma_{1})$ and include only one copy of the fermion multiplet $\Sigma_{R}$.
If the fermions are heavier than the scalars ($M_{\Sigma}>M_{\Phi}$), the scalars can only decay to the gauge bosons, which follows from Eqs. 3 and 4777For a detailed study of the case $M_{\Sigma}>M_{\Phi}$, please see [Kumar:2021umc]. In the converse case, the scalars can decay to the fermions as well, in addition to the gauge bosons. Thus, in the latter case, we observe two scenarios:
@@
Since we consider $M_{\Phi}>M_{\Sigma}$, the quintuplets do not decay to scalars. However, by virtue of their mixing with the SM leptons which arises due to the Yukawa interaction in Eq. 4 when $\phi^{0}$ obtains the VEV, they can decay directly to an SM fermion and a gauge boson. The mixing of the neutral fermions with SM fermions can be written as,
@@
Similarly, the mixing of the charged fermions with SM fermions can be termed as,
@@
Diagonalisation of the mass matrices leads to the following mixing matrices [Grimus:2000vj, Kumericki:2012bh],
@@
Here, $U_{PMNS}$ is the $3\times 3$ matrix that diagonalises the light neutrino masses and $V=\frac{v_{4}}{\sqrt{2}}y_{\nu}{M_{\Sigma}}^{-1}$. The value of $V$ is constrained from the neutrino masses and is estimated to be $\sim 3.5\times 10^{-7}$ [Kumericki:2012bh, Yu:2015pwa]. The physical states obtained after mixing are denoted by a $\sim$’ over the fields. For simplicity, however, we drop the $\sim$’ hereafter. The $U_{PMNS}$ matrix can be expressed as [ParticleDataGroup:2022pth]
@@
where $c_{ij}=\cos\theta_{ij}$, $s_{ij}=\sin\theta_{ij}$; $\theta_{12},\theta_{23}$ and $\theta_{13}$ are the light neutrino mixing angles; $\phi$ and $\phi^{\prime}$ are Majorana phases; $\delta$ is the Dirac CP violating phase. Data obtained from solar neutrino [SNO:2002tuh, Super-Kamiokande:2001ljr] and atmospheric neutrino searches [Super-Kamiokande:1998kpq] constraints the neutrino masses and the mixing involved. As summarized in [ParticleDataGroup:2022pth], the allowed values of mixing angles and phases are $\theta_{ij}\in[0,\pi/2]$, $\delta\in[0,\pi]$. Considering the large uncertainty in phase values and negligible impact of $\delta$ on determination of Yukawa couplings in seesaw mechanisms [Casas:2001sr], all phases are assumed to be zero for simplicity. Further, we consider normal hierarchy and use the following best-fit values for for $\theta_{ij}$[Esteban:2018azc]
@@
The gauge couplings of $\Sigma_{R}$ can be written as,
@@
We combine the similarly charged components to form Dirac fermions,
@@
$\mathcal{L}_{gauge,SM}$ $\supset$ $\frac{g}{c_{w}}\bigg[\overline{\nu}_{L}\gamma^{\mu}(\frac{1}{2})\nu_{L}+\overline{l}_{L}\gamma^{\mu}(s_{w}^{2}-\frac{1}{2})l_{L}\bigg]Z_{\mu}+\frac{g}{\sqrt{2}}\bigg[\overline{l}_{L}\gamma^{\mu}\nu_{L}W^{-}_{\mu}+h.c.\bigg],$ (13)
Substituting the mass eigenstates from Eq. 7 and Eq. 8 in Eq. 11 and Eq. 13 leads to the following interaction Lagrangian which describes the interaction among the quintuplet fermions, SM leptons and gauge bosons.
@@ 3.1 Production at the LHC
The charged scalars are produced in $pp$ collisions via SM gauge bosons in the $s$-channel. We take particular interest in the pair and associated productions of the charged scalars due to their large cross-sections compared to the production cross-sections (pair or associated) of the neutral scalars viz.,
$pp\rightarrow\phi^{++}\phi^{--};~~~~~pp\rightarrow\phi^{++}\phi_{2}^{-}+h.c.;$
$pp\rightarrow\phi_{1}^{+}\phi_{1}^{-};~~~~~~~~pp\rightarrow\phi_{2}^{+}\phi_{2}^{-}.$ (16)
We plot these cross-sections ($\sigma$) as a function of the scalar mass at 14 TeV LHC in Fig. 1. The production cross-section for $\phi_{1}^{+}\phi_{1}^{-}$ differs from the production cross-section for $\phi_{2}^{+}\phi_{2}^{-}$, and is larger by a factor of $\sim 10$.
Figure 1: (left) Cross-sections for the doubly charged scalar pair production and associated production, (right) pair production cross-section of the singly charged scalar as a function of their mass at 14 TeV LHC.
For the doubly charged scalar, the associated production ($\phi^{++}\phi_{2}^{-}+h.c.$) is larger than the pair production. There is no associated production of doubly charged scalar with $\phi_{1}$ because the gauge interaction in Eq. 2 does not couple $\phi^{\pm\pm}$ (isospin $3/2$) with $\phi_{1}^{\pm}$ (isospin 3/2-3/2). Further, the associated production takes place via $W$-boson as a mediator. Since $\sigma(pp\rightarrow W^{+})>\sigma(pp\rightarrow W^{-})$, the production cross-section of $\phi_{2}^{+}\phi^{--}$ is less than that of $\phi^{++}\phi_{2}^{-}$.
@@ 3.2 Fermiophobic and Fermiophilic Decays
When $M_{\Phi}<M_{\Sigma}$, the quadruplet scalars decay only to the gauge bosons and behave as purely fermiophobic [Nomura:2017abu, Kumar:2019tat, Kumar:2021umc]. The decay modes are
$\phi^{\pm}_{1,2}$ $\rightarrow$ $W^{\pm}Z$
$\phi^{\pm\pm}$ $\rightarrow$ $W^{\pm}W^{\pm}$
$\phi^{0}$ $\rightarrow$ $W^{+}W^{-}$ (17)
@@
$\phi_{1}^{\pm}$ $\rightarrow$ $\Sigma^{0}l^{\pm}/\Sigma^{\pm}\nu$
$\phi_{2}^{\pm}$ $\rightarrow$ $\Sigma^{\pm}\nu/\Sigma^{\pm\pm}l^{\mp}$
$\phi^{\pm\pm}$ $\rightarrow$ $\Sigma^{\pm\pm}\nu/\Sigma^{\pm}l^{\pm}$
$\phi^{0}$ $\rightarrow$ $\overline{\nu}_{l}\Sigma^{0}/\overline{l}\Sigma^{-}$ (18)
@@
In the fermiophilic decay modes, the quintuplet fermion further decay to SM gauge bosons and leptons via the following modes:
$\Sigma^{0}$ $\rightarrow$ $l^{-}W^{+},\nu_{l}Z$
$\Sigma^{\pm}$ $\rightarrow$ $l^{\pm}Z,\nu W^{\pm}$
$\Sigma^{\pm\pm}$ $\rightarrow$ $W^{\pm}l^{\pm}$ (19)

### v2

@@ Abstract
Beyond the Standard Model scenarios introduce additional scalar and fermion multiplets, which influence neutrino mass generation mechanisms and yield distinctive collider signatures. This work focuses on a particular scenario involving a fermion quintuplet and a scalar quadruplet. The study examines the production and decay of the scalar quadruplet components at the Large Hadron Collider (LHC), emphasizing how their decay patterns, fermiophobic versus fermiophilic, depend on mass differences and Yukawa couplings with the fermion multiplets. This study provides an overview of possible signals at the LHC, along with a detailed collider analysis focused on final states containing at least four leptons and two jets, in which the masses of the scalars and fermions are reconstructed successfully. Standard Model backgrounds are also incorporated in the study, with multivariate techniques leveraged via Boosted Decision Trees. Results indicate discovery potential for scalar masses around 600700600–700 GeV and exclusion sensitivity extending beyond $1$ TeV, highlighting the promising experimental signatures of the model and its role in probing new physics at colliders.
@@ 1 Introduction
The Standard Model (SM), while spectacularly successful, is fundamentally incomplete. Its inability to explain phenomena such as neutrino masses, the nature of dark matter, and the hierarchy problem, among other issues, strongly motivates physics beyond the SM (BSM). Many sophisticated models are proposed to overcome the shortcomings of SM [45, 57, 44, 70, 61, 39, 60, 17, 48, 14, 58, 64, 30]. In many of these models, extra scalar degrees of freedom are either necessary or appear as a consequence of the fundamental theory. However, simple extensions of SM such as Two Higgs Doublet Model (2HDM) [25], Type-II seesaw with scalar triplets [63], appear with additional scalar multiplets. In these models, tiny neutrino masses arise via the seesaw mechanism. However, tree-level seesaw models face tension between the seesaw scale and couplings [24, 33]. A resolution to this is to suppress the tree-level contribution and make the leading contribution come from loop-level diagrams of the dimension-5 operator [23, 55], or to raise the dimension of the effective operator[19, 53, 59, 40]. The appearance of large scalar multiplets is very natural in these types of models[42, 41]. Another alternative method to lower the seesaw scale is by the “cascade seesaw mechanism”[33], where the neutrino masses are generated at tree level via a combination of Type-I and Type-III seesaw mechanisms. These models include both scalar and fermion multiplets as part of their exotic particle content.
@@
We explore the production and decays of the quadruplet scalar (comprising neutral, singly charged, and doubly charged components) at the LHC. The decay modes of these quadruplet scalars depend critically on the mass difference between the scalar and fermion multiplets. The Yukawa coupling between the scalar and fermion multiplets also plays a major role in their decays. We show that there are two distinct decay patterns where the scalar can either decay to the gauge bosons directly (when the Yukawa coupling is small) or they first decay to quintuplet fermions which subsequently decay to SM particles (when the Yukawa coupling is large). These two scenarios are named fermiophobic and fermiophilic, respectively. We discuss the possible collider signatures in these two scenarios, along with the existing limits from the collider studies.
The introduction of two distinct $SU(2)$ multiplets as the BSM particle content poses significant challenges for the phenomenology. We discuss the pair and associated production of charged scalars within the mass range 3001000300-1000 GeV, at 14 TeV LHC. Beyond identifying channels with superior discovery potential, this paper focuses on those enabling mass reconstruction of exotic scalars and fermions—crucial for new particle discovery. Our detailed analysis of signal processes reveals that the final state with 4 or more leptons and 2 jets (fermiophilic scenario) offers sufficient sensitivity to probe both the quartet scalar and quintuplet fermions at the LHC. We analyze the most relevant SM background that exhibits overlapping kinematics, with a focus on the key reconstructed observables that offer some discrimination power. The sensitivity in these channels is estimated by performing a multivariate analysis (MVA) using the Boosted Decision Tree (BDT) framework. We also demonstrate the reconstruction of the quintuplet scalars and quintuplet fermion from their decay products.
@@ 2 Model
This work is motivated by the models studied previously in [53, 59, 50]. To keep the analysis minimal and transparent, we consider the simplest scenario where the BSM particle content comprises a quadruplet scalar ($\Phi_{4}=\left(\phi^{++},\phi^{+}_{2},\phi^{0},\phi^{-}_{1}\right)^{T}$) and a quintuplet fermion ($\Sigma_{R}=\left[\Sigma_{1}^{++},\Sigma_{1}^{+},{\Sigma^{0}},\Sigma_{2}^{-},\Sigma_{2}^{--}\right]_{R}^{T}$) only. We denote the scalar and the fermion masses by $M_{\Phi}$ and $M_{\Sigma}$, respectively  [^5] The mass of the scalar comes from the scalar potential $V=\mu^{2}H^{\dagger}H+M_{\Phi}^{2}\Phi_{4}^{\dagger}\Phi_{4}$..
The kinetic term of the quadruplet scalar, obtained by expanding the tensors, is
@@
where $m$ is the weak isospin and $s_{W}~(c_{W})=\sin\theta_{W}~(\cos\theta_{W})$, $\theta_{W}$ being the Weinberg angle. From here, we obtain the interactions responsible for the production of the scalars viz.,
@@
Here, the neutral component of ${\Phi}_{4}$ gets a non-zero VEV, $\frac{v_{4}}{\sqrt{2}}$, induced by virtue of its quartic coupling with the SM doublet Higgs. The VEV is constrained from the measurement of the $\rho$ parameter [53, 72, 59]. [^6] These models are further constrained from muon (g2)(g-2) and lepton flavor violation experiments which is discussed in [59, 50].. The decay of the scalars to the gauge bosons is governed by the following terms:
@@
It is important to note that the decay of the $\Phi_{4}$ to gauge bosons depends upon its VEV, $v_{4}$.
The Yukawa interaction between the quadruplet scalar and the quintuplet fermion can be expressed as
@@
This Yukawa interaction induces diagrams that lead to tiny neutrino masses [59, 50]. The induced VEV allows decent Yukawa coupling for quintuplet masses $\leq 1$ TeV as neutrino mass, $m_{\nu}\propto v_{4}^{2}~y_{\nu}M_{\Sigma}^{-1}y_{\nu}^{T}$ [59]. For example, we find that for small $y_{\nu}$ ($\sim 0.0007$), $v_{4}\sim$ 1 GeV and for large $y_{\nu}$ ($\sim 1$), $v_{4}\sim 0.0002$ GeV. In order to generate non-zero masses for all three SM neutrinos, we need three copies of the quintuplets [29, 62]. For the purpose of this study, we assume that $M(\Sigma_{2},~\Sigma_{3})>>M(\Sigma_{1})$ and include only one copy of the fermion multiplet $\Sigma_{R}$.
If the fermions are heavier than the scalars ($M_{\Sigma}>M_{\Phi}$), the scalars can only decay to the gauge bosons, which follows from Eqs. 3 and 4777For a detailed study of the case $M_{\Sigma}>M_{\Phi}$, please see [51]. In the converse case, the scalars can decay to the fermions as well, in addition to the gauge bosons. Thus, in the latter case, we observe two scenarios:
@@
Since we consider $M_{\Phi}>M_{\Sigma}$, the quintuplets do not decay to scalars. However, by virtue of their mixing with the SM leptons, which arises due to the Yukawa interaction in Eq. 4 when $\phi^{0}$ obtains the VEV, they can decay directly to an SM fermion and a gauge boson. The mixing of neutral fermions with SM fermions can be written as
@@
Similarly, the mixing of charged fermions with SM fermions can be written as
@@
Diagonalization of the mass matrices leads to the following mixing matrices [43, 53]
@@
Here, $U_{PMNS}$ is the $3\times 3$ matrix that diagonalizes the light neutrino masses and $V=\frac{v_{4}}{\sqrt{2}}y_{\nu}{M_{\Sigma}}^{-1}$. The value of $V$ is constrained by neutrino masses and is estimated to be $\sim 3.5\times 10^{-7}$ [53, 72]. The physical states obtained after mixing are denoted by a $\sim$’ over the fields. For simplicity, however, we drop the $\sim$’ hereafter. The $U_{PMNS}$ matrix can be expressed as [71]
@@
where $c_{ij}=\cos\theta_{ij}$, $s_{ij}=\sin\theta_{ij}$; $\theta_{12},\theta_{23}$ and $\theta_{13}$ are the light neutrino mixing angles; $\phi$ and $\phi^{\prime}$ are Majorana phases; $\delta$ is the Dirac CP violating phase. Data obtained from solar neutrino [15, 37] and atmospheric neutrino searches [38] constrain the neutrino masses and the mixing involved. As summarized in [71], the allowed values of mixing angles and phases are $\theta_{ij}\in[0,\pi/2]$, $\delta\in[0,\pi]$. Considering the large uncertainty in phase values and negligible impact of $\delta$ on determination of Yukawa couplings in seesaw mechanisms [29], all phases are assumed to be zero for simplicity. Further, we consider normal hierarchy and use the following best-fit values for $\theta_{ij}$[35]
@@
The gauge couplings of $\Sigma_{R}$ can be written as
@@
We combine the similarly charged components to form Dirac fermions
@@
$\mathcal{L}_{gauge,SM}$ $\supset$ $\frac{g}{c_{w}}\bigg[\overline{\nu}_{L}\gamma^{\mu}(\frac{1}{2})\nu_{L}+\overline{l}_{L}\gamma^{\mu}(s_{w}^{2}-\frac{1}{2})l_{L}\bigg]Z_{\mu}+\frac{g}{\sqrt{2}}\bigg[\overline{l}_{L}\gamma^{\mu}\nu_{L}W^{-}_{\mu}+h.c.\bigg].$ (13)
Substituting the mass eigenstates from Eqs. 7 and 8 in Eqs. 11 and 13 leads to the following interaction Lagrangian, which describes the interaction among the quintuplet fermions, SM leptons, and gauge bosons.
@@ 3.1 Production at the LHC
The charged scalars are produced in $pp$ collisions via SM gauge bosons in the $s$-channel. We take particular interest in the pair and associated productions of the charged scalars due to their large cross sections compared to the production cross sections (pair or associated) of the neutral scalars viz.,
$pp\rightarrow\phi^{++}\phi^{--},~~~~~pp\rightarrow\phi^{++}\phi_{2}^{-}+h.c.;$
$pp\rightarrow\phi_{1}^{+}\phi_{1}^{-},~~~~~~~~pp\rightarrow\phi_{2}^{+}\phi_{2}^{-}.$ (16)
We plot these cross sections ($\sigma$) as a function of the scalar mass at 14 TeV LHC in Fig. 1. The production cross section for $\phi_{1}^{+}\phi_{1}^{-}$ differs from the production cross section for $\phi_{2}^{+}\phi_{2}^{-}$, and is larger by a factor of $\sim 10$.
Figure 1: (left) Cross sections for the doubly charged scalar pair production and associated production, (right) pair production cross section of the singly charged scalar as a function of their mass at 14 TeV LHC.
For the doubly charged scalar, the associated production ($\phi^{++}\phi_{2}^{-}+h.c.$) is larger than the pair production. There is no associated production of doubly charged scalar with $\phi_{1}$ because the gauge interaction in Eq. 2 does not couple $\phi^{\pm\pm}$ (isospin $3/2$) with $\phi_{1}^{\pm}$ (isospin 3/2-3/2). Further, the associated production takes place via a $W$-boson as a mediator. Since $\sigma(pp\rightarrow W^{+})>\sigma(pp\rightarrow W^{-})$, the production cross section of $\phi_{2}^{+}\phi^{--}$ is less than that of $\phi^{++}\phi_{2}^{-}$.
@@ 3.2 Fermiophobic and Fermiophilic Decays
When $M_{\Phi}<M_{\Sigma}$, the quadruplet scalars decay only to the gauge bosons and behave as purely fermiophobic [59, 50, 51]. The decay modes are
$\phi^{\pm}_{1,2}$ $\rightarrow$ $W^{\pm}Z,$
$\phi^{\pm\pm}$ $\rightarrow$ $W^{\pm}W^{\pm},$
$\phi^{0}$ $\rightarrow$ $W^{+}W^{-}.$ (17)
@@
$\phi_{1}^{\pm}$ $\rightarrow$ $\Sigma^{0}l^{\pm}/\Sigma^{\pm}\nu,$
$\phi_{2}^{\pm}$ $\rightarrow$ $\Sigma^{\pm}\nu/\Sigma^{\pm\pm}l^{\mp},$
$\phi^{\pm\pm}$ $\rightarrow$ $\Sigma^{\pm\pm}\nu/\Sigma^{\pm}l^{\pm},$
$\phi^{0}$ $\rightarrow$ $\overline{\nu}_{l}\Sigma^{0}/\overline{l}\Sigma^{-}.$ (18)
@@
In the fermiophilic decay modes, the quintuplet fermions further decay to SM gauge bosons and leptons via the following modes:
$\Sigma^{0}$ $\rightarrow$ $l^{-}W^{+},\nu_{l}Z,$
$\Sigma^{\pm}$ $\rightarrow$ $l^{\pm}Z,\nu W^{\pm},$
$\Sigma^{\pm\pm}$ $\rightarrow$ $W^{\pm}l^{\pm}.$ (19)